OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 729 m3?

OpenStudy (pawanyadav):

Volume of cube=side*side*side

OpenStudy (pawanyadav):

Side=(volume) ^1/3 = (729)^1/3=9 m

OpenStudy (anonymous):

thanks I just forgot how to do it but this reminded me. :3 imma medal you for your awesomeness

OpenStudy (pawanyadav):

Welcome